OpenStudy (anonymous):

6 radians is the same as _____degrees. Round to the nearest hundredth of a degree.

OpenStudy (abhisar):

Hello @countryrockbabe !

OpenStudy (anonymous):

hello please help

OpenStudy (abhisar):

6 radians = \([\large \sf 6\times\frac{180}{\pi}]°\)

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

Got it ? What would be the answer ?

OpenStudy (abhisar):

@countryrockbabe you probably didn't notice, but you posted your \(\large\sf\color{blue}{\underline{\href{ /study#/groups/mathematics}{Mathematics}}}\) question in the Feedback section. Next time when you ask a question, check if you're in the right section :)